OpenStudy (anonymous):

find the x-intercept of the parabola with vertex (-3,-14) and y intercept (0,13) in (x1,y1) (x2,y2) form using - (y-k)=a(x-h)^2

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

Please see the attachment. This method works for whenever you have the vertex (h,k) and another point on the quadratic.

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

oh i see the instruction are not clear at all are they?

OpenStudy (anonymous):

are you also tabbeyx33 because this is exactly what is posted there as well. i would just try to enter \[y+14=3(x+3)^2\] because i think the instructions are left over from some other problem.

OpenStudy (anonymous):

Oops, didn't finish the problem off.

OpenStudy (anonymous):

no my name is priscila and i just need an asnswer to putt lol.

OpenStudy (anonymous):

i have never heard of \[(x_1,y_1), (x_2,y_2)\] form for a parabola so maybe if i watch jabberwock i will learn something

OpenStudy (anonymous):

oh lord it says "find the x intercept" ok fine. start with \[y=3(x+3)^2-14\], set y = 0 and solve!

OpenStudy (anonymous):

\[0=3(x+3)^2-14\] \[14=3(x+3)^2\] \[\frac{14}{3}=(x+3)^2\] \[x+3=\pm\sqrt{\frac{14}{3}}\] \[x=-3\pm\sqrt{\frac{14}{3}}\]

OpenStudy (anonymous):

a more traditional way to do this is write \[y=3(x+3)^2-14\] and expand. you get \[y=3(x^2+6x+9)-14\] \[y=3x^2+18x+27-14\] \[y=3x^2+18x + 13\] set \[0=3x^2+18x + 13\] an use the quadratic formula to get \[x=\frac{-9\pm\sqrt{42}}{3}\]

OpenStudy (anonymous):

so x intercepts are \[(\frac{-9-\sqrt{42}}{3},0)\] and \[(\frac{-9+\sqrt{42}}{3},0)\]

OpenStudy (anonymous):

that is the answer you need to put in and if you want a decimal use a calculator to get it

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

i can try again

OpenStudy (anonymous):

no is ok i put it in already & it was wrong :( but is ok lol. ^_^

OpenStudy (anonymous):

i got another one

OpenStudy (anonymous):

yes i am sure. pretty sure anyway

OpenStudy (anonymous):

fire away

OpenStudy (anonymous):

nevermind i closed my test like a dumbass -____

OpenStudy (anonymous):

on line class? dang

OpenStudy (anonymous):

i need help with word problems.

OpenStudy (anonymous):

ask

OpenStudy (anonymous):

here you go.

OpenStudy (anonymous):

can you see it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

did you attach it?